Let # (a,b,c) = a(b + c). What is the value of the expression #(#(0,2,3), #(5,4,7), #(3,4,2))?
Answer by venugopalramana(3286)

Let # (a,b,c) = a(b + c). What is the value of the expression #(#(0,2,3), #(5,4,7), #(3,4,2))?
JUST PUT VALUES OF A,B,C FROM THE GIVEN DATA IN THE FORMULA GIVEN TO GET THE ANSWER
FORMULA IS
# (a,b,c) = a(b + c).
SO #(0,2,3)=0(2+3)=0..SINCE A=0,B=2,C=3
SIMILARLY #(5,4,7)=5(4+7)=5*11=55
#(3,4,2)=3(4+2)=3*6=18
SO #(#(0,2,3), #(5,4,7), #(3,4,2)) = #(0,55,18)=0(55+18)=0